An investor owns 5 000 IBM shares, with a market value of R230 per share. If IBM has a 3-for-1 subdivision, how many shares will the investor now own?

1 667

1 666

15 000

20 000
Answer #1

Total market value = 5,000 * 230 = 1,150,000

Price per share after split = 230 / 3 = 76.6667

Number of shares = 1,150,000 / 76.6667

Number of shares = 15,000